The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs. Solve using algebraic equations.
Thank you.

Answer by elima(1433)

The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs. Solve using algebraic equations.
=============================================
(4x-8)+(4x-8)+x=29
9x-16=29
9x=45
x=5
=============
4(5)-8=20-8=12
:)